OpenStudy (anonymous):

-3x – y = 1 x + 6y = 11...what are the 2 answers?

OpenStudy (anonymous):

Are they two separate equations?

OpenStudy (anonymous):

yes formated like this (___,___)

OpenStudy (anonymous):

Is it suppose to be -3x-y=1 And. x+6y=11

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If it is how I showed it then you will need to multiply either the top or bottom by a number to cancel out one of the letters. Example multiply the entire bottom row by the and then add the two rows together.

OpenStudy (anonymous):

Sorry multiply bottom by three

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

3(-x+6y=11)

OpenStudy (anonymous):

but there is 2 seperate answers

OpenStudy (anonymous):

a= im hungry

OpenStudy (tyteen4a03):

-3x – y = 1 x + 6y = 11 can be formulated into these 2 simultaneous equations: -3x - y = 11 x + 6y = 11 There are two ways to do this, but I'll pick the easier one: substitution. Firstly, make either variable the subject of one equation. I'll pick y and the first equation. Add y to both sides then subtract 11 from both sides to get y = -3x - 11. Then substitute this newfound equation into the second equation: x + 6(-3x-11) = 11 Process 6(-3x-11) and get x - 18x - 66 = 11. Add 66 to both sides and simplify x to get -17x = 77. Now divide -17 on both sides to get x = -77/17. Substitute x = -77/17 into x + 6y = 11 and continue to find y.